OVERTONE SPECTROSCOPY IN TRANS-FORMIC ACID

M. FREYTES, D. HURTMANS, S. KASSI, J. LIEVIN, \underline{J. VANDER AUWERA}, M. HERMAN, Laboratoire de Chimie Physique Moléculaire C. P. 160/09, Université Libre de Bruxelles, 50 Avenue F. D. Roosevelt, B-1050 Brussels, Belgium; A. CAMPARGUE, Laboratoire de spectrométrie Physique (CNRS, UMR C5588), Université Joseph Fourier de Grenoble, B.P. 87, 38042 Saint-Martin d’Héres Cédex, France.

Vibrational assignments of fundamental, combination and overtone bands in the main isotopomer of gaseous trans-formic acid are reported from spectra either newly or previously ^{a} recorded using high-resolution Fourier transform spectroscopy and intracavity laser absorption spectroscopy. A total of 62 bands, with 32 newly reported ones, were observed from the lowest energy band ν_{7} at 626.16 cm^{−1} up to 4ν_{1} at 13284.1 cm^{−1}. Among these bands, 43 were firmly assigned, and 16 tentatively. Effective vibrational constants were obtained. The normal modes of vibrations were further characterized using ab initio calculations providing fundamental band intensities and normal mode nuclear displacements. The analysis of the rotational structure in the first CH stretch overtone band (2ν_{2}) and in the second OH stretch overtone band (3ν_{1}) was performed. Some rotational information could also be obtained for 3ν_{2} and two close-lying bands, extracted from strong overlapping formic acid dimer bands.

aD. Hurtmans, F. Herregodts, M. Herman, J. Liévin, A. A. Kachanov and A. Campargue, J. Chem. Phys. \underline{113}, 1535 (2000).